OpenStudy (anonymous):

multiply and express in lowest terms x^2+3x-10/6x^3 * 3x^2/(x-2)

OpenStudy (anonymous):

(x+5)/(2 x)

OpenStudy (anonymous):

wow great how did you got there please explain

OpenStudy (anonymous):

how you do this multiply all the numerators and then denominators?? what i'm i doing wrong

OpenStudy (anonymous):

(x^2+3x-10) = (x+5)(x-2) :- (x+5)(x-2) (3x^2) ---------- * ------ 6X^3 (x-2) the (x-2) in the numerator cancels out (x-2) for the denominator then (3x^2)/(6x^3) = 1/2x :- (x+5)/(2x)

OpenStudy (anonymous):

thank you...

OpenStudy (anonymous):

npz, good luck!